OpenStudy (anonymous):

A certain material decays at a rate of 0.92% per year. How much of 260 grams of the material will be left in 11 years?

OpenStudy (anonymous):

26.31 is what i get

OpenStudy (anonymous):

can you show me your work? My teacher is a stickler for that XP

OpenStudy (anonymous):

260 x (.92% x 11) But i am not sure if that is right, get it double checked

OpenStudy (anonymous):

yes, use the formula N(t)=N(0)e^-kt N((t)=260e^(-0.92/100)t at t=11 years therefore N(11)= 260^(-0.92/100)(11)=234.98 gr approx

OpenStudy (anonymous):

for decay formula use N(t)=N(0)e^-kt for growth formula use N(t)=N(0)e^kt

OpenStudy (anonymous):

any questions guys?

Directrix (directrix):

I put the following and got the same answer. My question is if the rate percentage of decay is to be changed from a percent to a real number before computation. I did that. N = 260 * e ^- ((.0092)(11)) N = 234.976

OpenStudy (anonymous):

yeheyyy correct well lets all have more fun lol.....

OpenStudy (anonymous):

:))